# The algebra $\mathcal{D}(W)$ via strong Darboux transformations

### v1

@@ 4. The algebra $\mathcal{D}(W)$ for a Laguerre example
The irreducible weight matrix $W(x)=e^{-x}x^{\alpha}\begin{pmatrix}x(1+a^{2}x)&&ax\\ ax&&1\end{pmatrix}$ is a strong Darboux transformation of the direct sum of scalar Laguerre weights $w(x)=\begin{pmatrix}e^{-x}x^{\alpha+1}&&0\\ 0&&e^{-x}x^{\alpha}\end{pmatrix}$.
@@
$D=-\partial^{2}xI-\partial\begin{pmatrix}\alpha+2-x&&0\\ 0&&\alpha+1-x\end{pmatrix}+\begin{pmatrix}1+\frac{1}{a^{2}}&&0\\ 0&&+\frac{1}{a^{2}}\end{pmatrix}\in\mathcal{D}(w)$
@@
$\mathcal{N}=W\mathcal{V}^{\ast}w^{-1}=\partial\begin{pmatrix}-ax&&x\\ -1&&0\end{pmatrix}+\begin{pmatrix}-a-\frac{1}{a}&&\alpha+1\\ 0&&\frac{1}{a}\end{pmatrix}.$
Now, we need to prove that $\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{w}$, for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C})[x]$. We observe that $W(x)=T(x)w(x)T(x)^{\ast}$ with $T(x)=\begin{pmatrix}1&&ax\\ 0&&1\end{pmatrix}$ and
@@
It is a matter of integration by parts, to obtain the differential operator $\mathcal{V}T=\partial\begin{pmatrix}0&&x\\ -1&&0\end{pmatrix}+\begin{pmatrix}-\frac{1}{a}&&-x+\alpha+1\\ 0&&\frac{1}{a}\end{pmatrix}$ is $w$-adjointable because $\lim_{x\to 0,\infty}P(x)\begin{pmatrix}0&&x\\ -1&&0\end{pmatrix}w(x)Q(x)^{\ast}=0$ for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C})[x]$. The $w$-adjoint of $\mathcal{V}T(x)$ is given by
$(\mathcal{V}T)^{\dagger}=w(\mathcal{V}T)^{\ast}w^{-1}=\partial\begin{pmatrix}0&&x\\ -1&&0\end{pmatrix}+\begin{pmatrix}-\frac{1}{a}&&\alpha+1\\ -1&&\frac{1}{a}\end{pmatrix}.$
@@
$\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P\cdot(\mathcal{V}T),QT\rangle_{w}=\langle P,QT\cdot(\mathcal{V}T)^{\dagger}\rangle_{w}=\langle P,Q\cdot\mathcal{N}\rangle_{w}.$
Therefore, $W$ is a strong Darboux transformation of $w$. ∎
@@
The sequence of monic orthogonal polynomials with respect to $w(x)=\begin{pmatrix}e^{-x}x^{\alpha+1}&&0\\ 0&&e^{-x}x^{\alpha}\end{pmatrix}$ is given by $P_{n}(x)=\operatorname{diag}(L_{n}^{\alpha+1}(x),L_{n}^{\alpha}(x))$. Hence, from Proposition 3.2, $Q_{n}(x)=P_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $W$. ∎
In order to determine the algebra $\mathcal{D}(W)$, we begin by studying some properties of the algebra $\mathcal{D}(w)$, for $w(x)=e^{-x}x^{\alpha+1}\oplus e^{-x}x^{\alpha}$. We know (see [3]) that
(11) $\mathcal{D}(w)=\left\{\begin{pmatrix}p_{1}(\delta_{\alpha+1})&&\mathcal{N}p_{2}(\delta_{\alpha})\\ \mathcal{V}p_{3}(\delta_{\alpha+1})&&p_{4}(\delta_{\alpha})\end{pmatrix}\biggm{|}p_{1},\,p_{2},\,p_{3},\,p_{4}\in\mathbb{C}[x]\right\},$
@@
The $w$-symmetric differential operators in $\mathcal{D}(w)$ satisfy the following property.
@@
Let $D=\sum_{j=0}^{n}\partial^{j}F_{j}\in\mathcal{S}(w)$ be a $w$-symmetric differential operator or order $n$. If $n$ is even, $n=2m$ for some $m\geq 0$, then the leading coefficient of $D$ satisfies
@@
Let $D$ be a $w$-symmetric operator of order $n$. Since $D=D^{\dagger}$, from (6) with $k=n$, we get that the leading coefficient of $D$ satisfies $F_{n}(x)w(x)=(-1)^{n}w(x)F_{n}(x)^{\ast}$. Now the statement follows easily from (11). ∎
@@
The weight $W(x)$ is a strong Darboux transformation of $w(x)=e^{-x}x^{\alpha+1}\oplus e^{-x}x^{\alpha}$ with strong Darboux transformer
@@
Let us consider the $w$-symmetric differential operators
@@
Let $E=\sum_{j=0}^{n}\partial^{j}F_{j}(x)$ be a differential operator in $\mathcal{S}(W)$ of order $n$. By Theorem 3.9 we have that $\mathcal{V}E\mathcal{N}\in\mathcal{S}(w)$.
@@
Now we will prove that the algebra $\mathcal{D}(W)$ is generated by =$\mathcal{M}=\{I,D_{1}.D_{2},D_{3},D_{4},D_{5}\}$. Let $E\in\mathcal{S}(W)$ be a differential operator of order $n$. For $n=0$, we get that $E=k_{1}I$ for some $k_{1}\in\mathbb{R}$. Also, it is easy to see that there are no differential operators of order one in $\mathcal{D}(W)$. We proceed by induction on $n$ and consider $n>1$.
For $n=2m$, using (12), we have that the leading coefficient of $E$ is a linear combination of the leading coefficients of $D_{2}$ and $D_{4}$ multiplied by $x^{m-1}$. Thus the differential operator $\mathcal{A}=E-D_{1}^{m-1}\big{(}k_{1}D_{2}+k_{4}D_{4}\big{)}$ has order less than $n$, hence by the inductive hypothesis, the operator $E$ is generated by $\mathcal{M}$. For $n=2m+1$, the leading coefficient of $E$ is a linear combination of the leading coefficients of $D_{3}$ and $D_{5}$ multiplied by $x^{m+1}$. See (13). Thus the differential operator $E-D_{1}^{m-1}\big{(}k_{2}D_{3}+k_{3}D_{5}\big{)}$ is of order less than $12m+$ 1 and by inductive hypothesis it follows that $E$ is generated by $\mathcal{M}$. Thus, $\mathcal{D}(W)$ is generated by $\{I,D_{1},D_{2},D_{3},D_{4},D_{5}\}$.
@@ 5. The algebra $\mathcal{D}(W)$ for Hermite examples
There are only three families $2\times 2$ of irreducible weight matrices of Hermite-type such that their algebra $\mathcal{D}(W)$ admits a second-order differential operator with leading coefficient $F_{2}(x)=I$. See [10]. They are
@@
(31) $W(x)$ $=e^{-x^{2}}\begin{pmatrix}1+a^{2}x^{2}&&ax\\ ax&&1\end{pmatrix}=\begin{pmatrix}1&&ax\\ 0&&1\end{pmatrix}\begin{pmatrix}e^{-x^{2}}&&0\\ 0&&e^{-x^{2}}\end{pmatrix}\begin{pmatrix}1&&0\\ ax&&1\end{pmatrix}$
@@
This section aims to determine the algebra $\mathcal{D}(W)$ for the weights given in (22) and (31), taking full advantage of the fact that the weight $W$ can be obtained as a Darboux transformation of the scalar weight $w(x)=e^{-x^{2}}I$.
@@
If fact, let $w_{1},w_{2}$ be scalar weights supported on $\mathbb{R}$. If $W$ is a Darboux transformation of $w(x)=\begin{pmatrix}w_{1}(x)&&0\\ 0&&w_{2}(x)\end{pmatrix}$ then there exist degree-preserving differential operators $\mathcal{V}$ and $\mathcal{N}$ as in the Definition 3.1. We observe that the 00-th order differential operator $\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$ belongs to $\mathcal{D}(w)$. By Proposition 3.3, we have that the differential operator $\mathcal{N}\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}\mathcal{V}\in\mathcal{D}(W)$. However, its leading coefficient is singular, which implies that it is not a polynomial in $D$. This gives a contradiction.
Let us consider a direct sum of scalar Hermite weights, $w(x)=e^{-x^{2}}I$. From [3], we know the structure of the algebra of this weight $w(x)$
(41) $\mathcal{D}(w)=\left\{\begin{pmatrix}p_{1}(\delta)&&p_{2}(\delta)\\ p_{3}(\delta)&&p_{4}(\delta)\end{pmatrix}\biggm{|}p_{1},\,p_{2},\,p_{3},\,p_{4}\in\mathbb{C}[x]\right\},$
where $\delta=\partial^{2}-2\partial x$ is the classical (scalar) Hermite differential operator. In particular, $\mathcal{D}(w)$ is a non-commutative algebra, and every differential operator within it is of even order.
@@
The leading coefficient of a $w$-symmetric differential operator $D=\displaystyle\sum_{j=0}^{2m}\partial^{j}F_{j}$ is
@@
From the symmetry equations given in Theorem 2.5, we obtain that the leading coefficient of $D$ satisfies $F_{2m}(x)w(x)=w(x)F_{2m}(x)^{\ast}$. Now, the statement follows easily from (41). ∎
@@ 5.1. Example I
First of all, we will prove here that $W$ is a Darboux transformation of the scalar weight $w(x)=e^{-x^{2}}$. Consequently, we will provide an explicit expression for a sequence of orthogonal polynomials along with their three-term recurrence relation. Additionally, we will determine the generators of $\mathcal{D}(W)$, establish some algebraic relations among them, and identify the generators of the center $\mathcal{Z}(W)$ within $\mathcal{D}(W)$.
@@
The weight matrix $W(x)$ is a strong Darboux transformation of the direct sum of classical scalar Hermite weights $w(x)=e^{-x^{2}}I$.
@@
$D=\begin{pmatrix}\delta^{2}-6\delta+8+\frac{16}{a^{2}}&&0\\ 0&&\delta^{2}+2\delta+\frac{16}{a^{2}}\end{pmatrix}\in\mathcal{D}(w)$
@@
(42) $\begin{split}\mathcal{V}&=\partial^{2}\begin{pmatrix}0&&1\\ -1&&ax^{2}\end{pmatrix}+\partial\begin{pmatrix}0&&-4x\\ 0&&0\end{pmatrix}+\begin{pmatrix}\frac{4}{a}&&-2\\ 0&&\frac{4}{a}\end{pmatrix},\\ \mathcal{N}=W\mathcal{V}^{\ast}w^{-1}&=\partial^{2}\begin{pmatrix}ax^{2}&&-1\\ 1&&0\end{pmatrix}+\partial\begin{pmatrix}4ax&&4x\\ 0&&0\end{pmatrix}+\begin{pmatrix}2a+\frac{4}{a}&&2\\ 0&&\frac{4}{a}\end{pmatrix}.\end{split}$
@@
$\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{w}\qquad\text{for all }P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C})[x].$
@@
It is worth noting that the Darboux transformation between $W$ and $w$ cannot arise from the factorization of a second-order differential operator $D\in\mathcal{D}(w)$. Otherwise, $\mathcal{V}$ and $\mathcal{N}$ would be of order $1$. Then, by Proposition 3.3, the algebra $\mathcal{D}(W)$ would have the following second-order operator given by $\mathcal{N}\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}\mathcal{V}$, but its leading coefficient is a singular matrix. We will see later that the leading coefficient of the differential operators of order less than or equal to $2$ in the algebra $\mathcal{D}(W)$ is $\alpha I$ for some $\alpha\in\mathbb{C}$. Thus, it leads us to a contradiction.
@@
The monic orthogonal polynomials with respect to $w(x)=e^{-x^{2}}I$ is given by $P_{n}(x)=\operatorname{diag}(H_{n}(x),H_{n}(x))$. Since $W(x)$ is a strong Darboux transformation of $w(x)$, from Proposition 3.2, we obtain that $Q_{n}(x)=P_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $W$. ∎
@@
Now, we aim to determine the algebra $\mathcal{D}(W)$. From Corollary 3.11, we have that all differential operator $D\in\mathcal{D}(W)$ has even order. We use the symmetry equations (8) to obtain that the algebra $\mathcal{D}(W)$ contains the second-order $W$-symmetric differential operator
@@
where $E_{1}=\begin{pmatrix}0&&1\\ 1&&0\end{pmatrix}$, $E_{2}=\begin{pmatrix}0&&0\\ 0&&1\end{pmatrix}$, $E_{3}=\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$ and $E_{4}=\begin{pmatrix}0&&i\\ -i&&0\end{pmatrix}$ which are $w$-symmetric differential operators.
@@
For any $W$-symmetric differential operator $D=\sum_{j=0}^{2n}\partial^{j}F_{j}$ in $\mathcal{D}(W)$ we have that $\mathcal{V}D\mathcal{N}\in\mathcal{S}(w)$, where $\mathcal{V}$ and $\mathcal{N}$ are the differential operators in (42). The leading coefficient of $\mathcal{V}$ is $\begin{pmatrix}0&&1\\ -1&&ax^{2}\end{pmatrix}$ and the leading coefficient of $\mathcal{N}$ is $\begin{pmatrix}ax^{2}&&-1\\ 1&&0\end{pmatrix}$ and the leading coefficient of $\mathcal{V}D\mathcal{N}$ is $\begin{pmatrix}k_{1}&&k_{2}+ik_{3}\\ k_{2}-ik_{3}&&k_{4}\end{pmatrix}$ for some $k_{1},k_{2},k_{3},k_{4}\in\mathbb{R}$. Thus we obtain that the leading coefficient of $D$ is
@@
Hence, from Proposition 2.2, $D_{4}$ and $D_{5}$ are generated by $\{I,D_{1},D_{2},D_{3}\}$ and this completes the proof. ∎
@@
Given a differential operator $D\in\mathcal{Z}(W)$, it follows that its adjoint $D^{\dagger}$ belongs to $\mathcal{Z}(W)$ because ${}^{\dagger}$ is an involution. Thus, it is easy to prove that $\mathcal{Z}(W)=\left(\mathcal{Z}(W)\cap\mathcal{S}(W)\right)\oplus i\left(\mathcal{Z}(W)\cap\mathcal{S}(W)\right)$. Then, it is sufficient to prove the statement for the symmetric operators in the center of $\mathcal{D}(W)$.
@@
Therefore, $Z_{1},Z_{2}$, and $Z_{3}$ are in $\mathcal{Z}(W)$. We also have that the leading coefficient of $Z_{1},Z_{2}$, and $Z_{3}$ is the identity $I$.

### v2

@@ 4. The algebra $\mathcal{D}(W)$ for a Laguerre example
The irreducible weight matrix $W(x)=e^{-x}x^{\alpha}\begin{pmatrix}x(1+a^{2}x)&&ax\\ ax&&1\end{pmatrix}$ is a strong Darboux transformation of the direct sum of scalar Laguerre weights $\widetilde{W}(x)=\begin{pmatrix}e^{-x}x^{\alpha+1}&&0\\ 0&&e^{-x}x^{\alpha}\end{pmatrix}$.
@@
$D=-\partial^{2}xI-\partial\begin{pmatrix}\alpha+2-x&&0\\ 0&&\alpha+1-x\end{pmatrix}+\begin{pmatrix}1+\frac{1}{a^{2}}&&0\\ 0&&\frac{1}{a^{2}}\end{pmatrix}\in\mathcal{D}(\widetilde{W})$
@@
$\mathcal{N}=W\mathcal{V}^{\ast}\widetilde{W}^{-1}=\partial\begin{pmatrix}-ax&&x\\ -1&&0\end{pmatrix}+\begin{pmatrix}-a-\frac{1}{a}&&\alpha+1\\ 0&&\frac{1}{a}\end{pmatrix}.$
Now, we need to prove that $\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{\widetilde{W}}$, for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C}[x])$. We observe that $W(x)=T(x)\widetilde{W}(x)T(x)^{\ast}$ with $T(x)=\begin{pmatrix}1&&ax\\ 0&&1\end{pmatrix}$ and
@@
It is a matter of integration by parts, to obtain the differential operator $\mathcal{V}T=\partial\begin{pmatrix}0&&x\\ -1&&0\end{pmatrix}+\begin{pmatrix}-\frac{1}{a}&&-x+\alpha+1\\ 0&&\frac{1}{a}\end{pmatrix}$ is $\widetilde{W}$-adjointable because $\lim_{x\to 0,\infty}P(x)\begin{pmatrix}0&&x\\ -1&&0\end{pmatrix}\widetilde{W}(x)Q(x)^{\ast}=0$ for all $P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C}[x])$. The $\widetilde{W}$-adjoint of $\mathcal{V}T(x)$ is given by
$(\mathcal{V}T)^{\dagger}=\widetilde{W}(\mathcal{V}T)^{\ast}\widetilde{W}^{-1}=\partial\begin{pmatrix}0&&x\\ -1&&0\end{pmatrix}+\begin{pmatrix}-\frac{1}{a}&&\alpha+1-x\\ 0&&\frac{1}{a}\end{pmatrix}.$
@@
$\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P\cdot(\mathcal{V}T),QT\rangle_{\widetilde{W}}=\langle P,QT\cdot(\mathcal{V}T)^{\dagger}\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{\widetilde{W}}.$
Therefore, $W$ is a strong Darboux transformation of $\widetilde{W}$. ∎
@@
The sequence of monic orthogonal polynomials with respect to $\widetilde{W}(x)=\begin{pmatrix}e^{-x}x^{\alpha+1}&&0\\ 0&&e^{-x}x^{\alpha}\end{pmatrix}$ is given by $P_{n}(x)=\operatorname{diag}(L_{n}^{\alpha+1}(x),L_{n}^{\alpha}(x))$. Hence, from Proposition 3.2, $Q_{n}(x)=P_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $W$. ∎
In order to determine the algebra $\mathcal{D}(W)$, we begin by studying some properties of the algebra $\mathcal{D}(\widetilde{W})$, for $\widetilde{W}(x)=e^{-x}x^{\alpha+1}\oplus e^{-x}x^{\alpha}$. We know (see [3]) that
(11) $\mathcal{D}(\widetilde{W})=\left\{\begin{pmatrix}p_{1}(\delta_{\alpha+1})&&\mathcal{N}p_{2}(\delta_{\alpha})\\ \mathcal{V}p_{3}(\delta_{\alpha+1})&&p_{4}(\delta_{\alpha})\end{pmatrix}\biggm{|}p_{1},\,p_{2},\,p_{3},\,p_{4}\in\mathbb{C}[x]\right\},$
@@
The $\widetilde{W}$-symmetric differential operators in $\mathcal{D}(\widetilde{W})$ satisfy the following property.
@@
Let $D=\sum_{j=0}^{n}\partial^{j}F_{j}\in\mathcal{S}(\widetilde{W})$ be a $\widetilde{W}$-symmetric differential operator or order $n$. If $n$ is even, $n=2m$ for some $m\geq 0$, then the leading coefficient of $D$ satisfies
@@
Let $D$ be a $\widetilde{W}$-symmetric operator of order $n$. Since $D=D^{\dagger}$, from (6) with $k=n$, we get that the leading coefficient of $D$ satisfies $F_{n}(x)\widetilde{W}(x)=(-1)^{n}\widetilde{W}(x)F_{n}(x)^{\ast}$. Now the statement follows easily from (11). ∎
@@
The weight $W(x)$ is a strong Darboux transformation of $\widetilde{W}(x)=e^{-x}x^{\alpha+1}\oplus e^{-x}x^{\alpha}$ with strong Darboux transformer
@@
Let us consider the $\widetilde{W}$-symmetric differential operators
@@
Let $E=\sum_{j=0}^{n}\partial^{j}F_{j}(x)$ be a differential operator in $\mathcal{S}(W)$ of order $n$. By Theorem 3.9 we have that $\mathcal{V}E\mathcal{N}\in\mathcal{S}(\widetilde{W})$.
@@
Now we will prove that the algebra $\mathcal{D}(W)$ is generated by =$\mathcal{M}=\{I,D_{1}.D_{2},D_{3},D_{4},D_{5}\}$. Let $E\in\mathcal{S}(W)$ be a differential operator of order $n$. For $n=0$, we get that $E=k_{1}I$ for some $k_{1}\in\mathbb{R}$.
There are no first-order differential operators in $\mathcal{D}(W)$, as none satisfy the symmetry equations given in (8). We proceed by induction on $n$ and consider $n>1$.
For $n=2m$, using (12), we have that the leading coefficient of $E$ is a linear combination of the leading coefficients of $D_{2}$ and $D_{4}$ multiplied by $x^{m-1}$. Thus the differential operator $\mathcal{A}=E-D_{1}^{m-1}\big{(}k_{1}D_{2}+k_{4}D_{4}\big{)}$ has order less than $n$, hence by the inductive hypothesis, the operator $E$ is generated by $\mathcal{M}$. For $n=2m+1$, the leading coefficient of $E$ is a linear combination of the leading coefficients of $D_{3}$ and $D_{5}$ multiplied by $x^{m+1}$, see (13). Thus the differential operator $E-D_{1}^{m-1}\big{(}k_{2}D_{3}+k_{3}D_{5}\big{)}$ is of order less than $12m+$ 1 and by inductive hypothesis it follows that $E$ is generated by $\mathcal{M}$. Thus, $\mathcal{D}(W)$ is generated by $\{I,D_{1},D_{2},D_{3},D_{4},D_{5}\}$.
@@ 5. The algebra $\mathcal{D}(W)$ for Hermite examples
There are only three families $2\times 2$ of irreducible weight matrices of Hermite-type such that their algebra $\mathcal{D}(W)$ admits a $W$-symmetric second-order differential operator with leading coefficient $F_{2}(x)=I$, see [10, Theorem 3.1]. They are
@@
(15) $W(x)$ $=e^{-x^{2}}\begin{pmatrix}1+a^{2}x^{2}&&ax\\ ax&&1\end{pmatrix}=\begin{pmatrix}1&&ax\\ 0&&1\end{pmatrix}\begin{pmatrix}e^{-x^{2}}&&0\\ 0&&e^{-x^{2}}\end{pmatrix}\begin{pmatrix}1&&0\\ ax&&1\end{pmatrix},$
@@
This section aims to determine the algebra $\mathcal{D}(W)$ for the weights given in (14) and (15), taking full advantage of the fact that the weight $W$ can be obtained as a Darboux transformation of the scalar weight $\widetilde{W}(x)=e^{-x^{2}}I$.
@@
If fact, let $w_{1},w_{2}$ be scalar weights supported on $\mathbb{R}$. If $W$ is a Darboux transformation of $\widetilde{W}(x)=\begin{pmatrix}w_{1}(x)&&0\\ 0&&w_{2}(x)\end{pmatrix}$ then there exist degree-preserving differential operators $\mathcal{V}$ and $\mathcal{N}$ as in the Definition 3.1. We observe that the 00-th order differential operator $\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$ belongs to $\mathcal{D}(\widetilde{W})$. By Proposition 3.3, we have that the differential operator $\mathcal{N}\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}\mathcal{V}\in\mathcal{D}(W)$. However, its leading coefficient is singular, which implies that it is not a polynomial in $D$. This gives a contradiction.
Let us consider a direct sum of scalar Hermite weights, $\widetilde{W}(x)=e^{-x^{2}}I$. From [3], we know the structure of the algebra of this weight $\widetilde{W}$
(17) $\mathcal{D}(\widetilde{W})=\left\{\begin{pmatrix}p_{1}(\delta)&&p_{2}(\delta)\\ p_{3}(\delta)&&p_{4}(\delta)\end{pmatrix}\biggm{|}p_{1},\,p_{2},\,p_{3},\,p_{4}\in\mathbb{C}[x]\right\},$
where $\delta=\partial^{2}-2\partial x$ is the classical (scalar) Hermite differential operator. In particular, $\mathcal{D}(\widetilde{W})$ is a non-commutative algebra, and every differential operator within it is of even order.
@@
The leading coefficient of a $\widetilde{W}$-symmetric differential operator $D=\displaystyle\sum_{j=0}^{2m}\partial^{j}F_{j}$ is
@@
From the symmetry equations given in Theorem 2.4, we obtain that the leading coefficient of $D$ satisfies $F_{2m}(x)\widetilde{W}(x)=\widetilde{W}(x)F_{2m}(x)^{\ast}$. Now, the statement follows easily from (17). ∎
@@ 5.1. Example I
First of all, we will prove here that $W$ is a Darboux transformation of the diagonal weight $\widetilde{W}(x)=e^{-x^{2}}$. Consequently, we will provide an explicit expression for a sequence of orthogonal polynomials along with their three-term recurrence relation. Additionally, we will determine the generators of $\mathcal{D}(W)$, establish some algebraic relations among them, and identify the generators of the center $\mathcal{Z}(W)$ within $\mathcal{D}(W)$.
@@
The weight matrix $W(x)$ is a strong Darboux transformation of the direct sum of classical scalar Hermite weights $\widetilde{W}(x)=e^{-x^{2}}I$.
@@
$D=\begin{pmatrix}\delta^{2}-6\delta+8+\frac{16}{a^{2}}&&0\\ 0&&\delta^{2}+2\delta+\frac{16}{a^{2}}\end{pmatrix}\in\mathcal{D}(\widetilde{W})$
@@
(18) $\begin{split}\mathcal{V}&=\partial^{2}\begin{pmatrix}0&&1\\ -1&&ax^{2}\end{pmatrix}+\partial\begin{pmatrix}0&&-4x\\ 0&&0\end{pmatrix}+\begin{pmatrix}\frac{4}{a}&&-2\\ 0&&\frac{4}{a}\end{pmatrix},\\ \mathcal{N}=W\mathcal{V}^{\ast}\widetilde{W}^{-1}&=\partial^{2}\begin{pmatrix}ax^{2}&&-1\\ 1&&0\end{pmatrix}+\partial\begin{pmatrix}4ax&&4x\\ 0&&0\end{pmatrix}+\begin{pmatrix}2a+\frac{4}{a}&&2\\ 0&&\frac{4}{a}\end{pmatrix}.\end{split}$
@@
$\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{\widetilde{W}}\qquad\text{for all }P,\,Q\in\operatorname{Mat}_{2}(\mathbb{C}[x]).$
@@
It is worth noting that the Darboux transformation between $W$ and $\widetilde{W}$ cannot arise from the factorization of a second-order differential operator $D\in\mathcal{D}(\widetilde{W})$. Otherwise, $\mathcal{V}$ and $\mathcal{N}$ would be of order $1$. Then, by Proposition 3.3, the algebra $\mathcal{D}(W)$ would have the following second-order operator given by $\mathcal{N}\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}\mathcal{V}$, but its leading coefficient is a singular matrix. We will see later that the leading coefficient of the differential operators of order less than or equal to $2$ in the algebra $\mathcal{D}(W)$ is $\alpha I$ for some $\alpha\in\mathbb{C}$. Thus, it leads us to a contradiction.
@@
The monic orthogonal polynomials with respect to $\widetilde{W}(x)=e^{-x^{2}}I$ are given by $P_{n}(x)=\operatorname{diag}(H_{n}(x),H_{n}(x))$. Since $W(x)$ is a strong Darboux transformation of $\widetilde{W}(x)$, from Proposition 3.2, we obtain that $Q_{n}(x)=P_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $W$. ∎
@@
Now, we aim to determine the algebra $\mathcal{D}(W)$. From Corollary 3.11, we have that every differential operator $D\in\mathcal{D}(W)$ has even order. We use the symmetry equations (8) to obtain that the algebra $\mathcal{D}(W)$ contains the second-order $W$-symmetric differential operator
@@
where $E_{1}=\begin{pmatrix}0&&1\\ 1&&0\end{pmatrix}$, $E_{2}=\begin{pmatrix}0&&0\\ 0&&1\end{pmatrix}$, $E_{3}=\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$ and $E_{4}=\begin{pmatrix}0&&i\\ -i&&0\end{pmatrix}$ which are $\widetilde{W}$-symmetric differential operators.
@@
For any $W$-symmetric differential operator $D=\sum_{j=0}^{2n}\partial^{j}F_{j}$ in $\mathcal{D}(W)$ we have that $\mathcal{V}D\mathcal{N}\in\mathcal{S}(\widetilde{W})$, where $\mathcal{V}$ and $\mathcal{N}$ are the differential operators in (18). The leading coefficient of $\mathcal{V}$ is $\begin{pmatrix}0&&1\\ -1&&ax^{2}\end{pmatrix}$, of $\mathcal{N}$ is $\begin{pmatrix}ax^{2}&&-1\\ 1&&0\end{pmatrix}$ and the leading coefficient of $\mathcal{V}D\mathcal{N}$ is $\begin{pmatrix}k_{1}&&k_{2}+ik_{3}\\ k_{2}-ik_{3}&&k_{4}\end{pmatrix}$, for some $k_{1},k_{2},k_{3},k_{4}\in\mathbb{R}$. Thus we obtain that the leading coefficient of $D$ is
@@
Hence, from Proposition 2.2, $D_{4}$ and $D_{5}$ are generated by $\{I,D_{1},D_{2},D_{3}\}$. ∎
@@
[…]
@@
For $D\in\mathcal{D}(W)$ we have $\Lambda_{n}(DZ_{i})=\Lambda_{n}(D)\Lambda_{n}(Z_{i})=\Lambda_{n}(Z_{i})\Lambda_{n}(D)=\Lambda_{n}(Z_{i}D)$ for all $n\geq 0$. By Proposition 2.2, it follows that $DZ_{i}=Z_{i}D$. Consequently, $Z_{1},Z_{2},Z_{3}$ belong to $\mathcal{Z}(W)$.